Stacy and Michael plan to retire in 40 years. How much do they need to deposit each month in a sinking fund in order to have $750,000 when they retire if they earn 4.8% compounded monthly. Please show work. Thanks you
Answer #1

(1+r) - 1 FutureValue = Periodic Payment *

where r is the rate of Interest for one compounding period i.e. 4.8%/ 12 = 0.4% OR 0.004

n is the no of compounding periods i.e. 40 year * 12 = 480 periods OR months

750,000 = Periodic Payment * (1 + 0.004) 80 - 1 0.0046.79488558522 - 1 750,000 = Periodic Payment * 0.004

750,000 = Periodic Payment * 1448.7213963

Periodic Payment = 750,000 / 1448.7213963

Periodic Payment = 517.70 OR $ 518

He needs to Deposit 517.70 OR $ 518 every month to get 750,000 at the end of 40 years.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.